Question 7 - Solve the following case (pg. 300 of the text book).

(5 marks)

Acadian Bakers

Acadian Bakers is a small but well known bakery, making and selling cakes and croissants in Houston. Its major customers, their locations on a coordinate system, and their weekly purchases are listed below. Acadian’s coordinates are (27, 36).

Coordinates

Customer

X

Y

Weekly Purchases ($)

B. Catering

22

31

80

B. Bar & Grill

38

38

70

H’s

38

36

90

H Club

38

37

120

H.C. Club

30

33

90

J’s Hideaway

−10

12

60

M’s

38

36

140

M. Club

−10

44

140

N.M.

47

33

300

K.S. Catering

33

33

50

Is Acadian centrally located to serve its customers?
Answer #1
Customer X Y Weekly Purchases ($) (l) lx ly
B. Catering 22 31 80 1760 2480
B. Bar & Grill 38 38 70 2660 2660
H’s 38 36 90 3420 3240
H Club 38 37 120 4560 4440
H.C. Club 30 33 90 2700 2970
J’s Hideaway -10 12 60 -600 720
M’s 38 36 140 5320 5040
M. Club -10 44 140 -1400 6160
N.M. 47 33 300 14100 9900
K.S. Catering 33 33 50 1650 1650
Total 1140 34170 39260

Center of gravity: x-coordinate = sum (lx)/sum(l) = 34170/1140= 29.97

y-coordinate =sum (ly)/sum(l) = 39260/1140 = 34.44

So Acadian’s coordinates should have been (29.97,34.44) to be located centrally which is not the case here so Acadian is not centrally located